OpenStudy (anonymous):

Write an equation for points (0,2) and (4,0). In point slope form, and slope intercept form.

OpenStudy (anonymous):

Use the formula Y2-Y1 / X2-X1 to find the slope of the equation

OpenStudy (anonymous):

I got -1/2 for the slope. Then what do I do?

OpenStudy (anonymous):

Ok that's right then use any one of the two points and make an equation in thif form y=mx+b

OpenStudy (anonymous):

so basically you just sub in the slope as ( m= -1/2)

OpenStudy (anonymous):

I got y= -1/2x+2

OpenStudy (anonymous):

and then use any two points to find out the b value in the equation above

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then you are done. Just need to rearrange the equation in point slope form which is y – y1 = m(x – x1)

OpenStudy (anonymous):

you can get more info on http://www.purplemath.com/modules/strtlneq2.htm